# Right-handed neutrino dark matter consistent with the generation of baryon number asymmetry

### v1

@@ 3 Consistency with leptogenesis
On the other hand, we find from the figure that $Y_{\eta_{R}}$ leaves $Y_{\eta_{R}}^{\rm eq}$ around $z_{2}=20$, which is considered to be caused due to the break of the balance between the $\eta_{R}$ production by the $N_{2}$ decay and its disappearance due to both the decay and the coannihilation. It shows that the late time additional $\eta_{R}$ production through the $N_{2}$ decay affects the estimation of the $N_{1}$ relic abundance. It forces us to change the strength of the $\eta_{R}$ decay and the coannihilation of $\eta_{R}$ in order to realize the required $N_{1}$ abundance since we impose simultaneous explanation of the DM abundance and the baryon number asymmetry on the model.eeeSince the $N_{2}$ decay is relevant to the leptogenesis, we do not change the parameters relevant to it here. In the figure, $Y_{\eta_{R}}$ and $Y_{N_{1}}$ are plotted by using blue and red lines, respectively. Different line types are used for different masses, $M_{1}$=1.5 TeV (solid), 150 GeV (dashed), and 1.5 GeV (dotted). $Y_{N_{1}}^{\infty}$ which can reproduce $\Omega_{N_{1}}h^{2}=0.12$ as the present DM abundance is shown by horizontal black solid, dashed and dotted lines for each $M_{1}$. Since $h_{1}$ does not affect the values of $h_{2,3}$ determined by the neutrino oscillation data as long as $h_{1}$ is small enough, it can be tuned to result in the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$. By tuning $h_{1}$ suitably to $h_{1}=2.3\times 10^{-9}$ and $2.3\times 10^{-8}$ for each case $M_{1}=$150 GeV and 1.5 GeV, $Y_{N_{1}}$ is found to converge to the required $Y_{N_{1}}^{\infty}$ at $z_{2}$ where $Y_{N_{1}}=Y_{N_{2}}$ is realized. These values of $h_{1}$ are much smaller than the ones used in Fig. 1. It is considered to be caused to cancel the effect of $\eta_{R}$ additionally produced through the $N_{2}$ decay by making the relative strength of the coannihilation of $\eta_{R}$ to the $\eta_{R}$ decay larger. The similar feature can be observed in the evolution of the number density of the keV scale $N_{1}$ also. For example, the DM relic abundance requires $h_{1}=6.5\times 10^{-5}$ for $M_{1}=5$ keV, and these parameters are confirmed not to affect the result of leptogenesis substantially. On the other hand, we find that the relation $Y_{N_{1}}+Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ cannot be changed by tuning the value of $h_{1}$ in the case $M_{1}=1.5$ TeV. This may be understood if we note that $Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ at $z_{2}=300$ where the coannihilation of $\eta_{R}$ is considered to be frozen out. It suggests that the freezeout of the coannihilation of $\eta_{R}$ plays an important role and then $\lambda_{3}$ and $\lambda_{4}$ have to be tuned especially when $M_{1}$ takes a similar order value to $M_{\eta_{R}}$.
@@
In the right panel, $Y_{N_{1}}$ and $Y_{\eta_{R}}$ are plotted for the case that $\lambda_{3}$ and $\lambda_{4}$ are shifted from the ones in the left panel to $(\lambda_{3},\lambda_{4})=(2.5,-1.0)$ keeping other parameters unchanged. While $Y_{N_{1}}$ in the case of $M_{1}=150$ GeV and 1.5 GeV is found to take a little bit smaller value than the one in the left panel, $Y_{N_{1}}+Y_{\eta_{R}}$ converges to $Y_{N_{1}}^{\infty}$ at $z_{2}=300$ in the case $M_{1}=1.5$ TeV. These examples show that $N_{1}$ in a wide mass range can explain the required relic abundance of DM in a consistent way with successful leptogenesis by tuning $h_{1}$ and $\lambda_{3,4}$ suitably. Couplings $\lambda_{3}$ and $\lambda_{4}$ play a crucial role especially when $M_{1}$ and $M_{\eta_{R}}$ take same order values since $Y_{N_{1}}^{\infty}~{}{{}^{>}_{\sim}}~{}Y_{\eta_{R}}^{\infty}$ should be satisfied in that case. We find that the lightest right-handed neutrino $N_{1}$ can be a good DM candidate consistently with leptogenesis and the model presents an interesting possibility such that three unsolved problems in the SM can be explained by low scale right-handed neutrinos simultaneously.
@@
Finally, we should address phenomenological signatures of the model. Extra charged and neutral scalars $\eta^{\pm}$ and $\eta_{R,I}$ might be directly examined in the future collider experiments [31]. However, if $N_{1}$ has a very small $h_{1}$, it might be difficult to distinguish the model from the one in which DM is $\eta_{R}$ through collider signatures. The existence of right-handed neutrinos and the additional inert doublet scalar at TeV scales are expected to cause a new non-negligible contribution at one-loop level to LFV processes like $\mu\rightarrow e\gamma$ and also the electric dipole moment of electron (EDME). It is useful to examine whether their present experimental bounds are satisfied and also to give their prospect in the future experiments. If we use the neutrino Yukawa couplings fixed by eqs. (7) and (12), we can estimate them definitely. Branching ratio of the LFV process $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ in the scotogenic model is expressed as [32]
@@
$B(\ell_{\alpha}\rightarrow\ell_{\beta}\gamma)=\frac{3\alpha}{64\pi(G_{F}M_{\eta}^{2})^{2}}\left|\sum_{k=1}^{3}U_{\alpha k}^{\ast}U_{\beta k}h_{k}^{2}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\right|^{2},$ (28)
@@
$d_{e}/e=\sum_{k=1}^{3}\frac{U_{ek}^{\ast}U_{ek}h_{k}^{2}m_{e}}{16\pi^{2}M_{\eta}^{2}}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\sin\gamma_{k},$ (29)
@@ 4 Summary
Both existences of neutrino mass and DM are unexplained problems in the SM. If they are solved in a common framework, it could be a promising candidate for new physics. The scotogenic model is such an example. The lightest right-handed neutrino is one of the DM candidates in this model. Although it could have a distinguishable character such that it cannot be found through direct detection experiments, it is known to be difficult to satisfy both constraints from the relic abundance and the LFV bounds simultaneously since they require incompatible magnitude for its Yukawa couplings. In order to escape this problem, additional interaction like $U(1)^{\prime}$ has been introduced to the right-handed neutrinos usually. Since it makes them interact with nuclei and then the direct detection experiments are relevant, the nature distinct from other ordinary WIMP candidates seems to be lost. In the present situation where the direct detection experiments exclude a lot of DM candidates, it may be an interesting possibility that DM is a right-handed neutrino which has no other interaction than the neutrino Yukawa couplings.
In this study, we propose a scenario in the original scotogenic model which allows the lightest right-handed neutrino can be DM without extra interaction. We show that the DM abundance and the baryon number asymmetry can be simultaneously explained if we assume certain mass degeneracy among new ingredients of the model. Although strict mass degeneracy is required, it is noticeable that three unsolved problems in the SM can be explained through the right-handed neutrinos simultaneously. An interesting point of this scenario is that the relic abundance of DM is determined not only by the interaction of DM itself but also by the interaction of its mother particle $\eta_{R}$. The model gives an example that the interaction searched by the direct detection experiments of DM is irrelevant to the interaction which determines the DM abundance.

### v2

@@ 3 Consistency with leptogenesis
In the figure, $Y_{\eta_{R}}$ and $Y_{N_{1}}$ are plotted by using blue and red lines, respectively. Different line types are used for different masses, $M_{1}$=1.5 TeV (solid), 150 GeV (dashed), and 1.5 GeV (dotted). A $Y_{N_{1}}^{\infty}$ value which can reproduce $\Omega_{N_{1}}h^{2}=0.12$ as the present DM abundance is shown by horizontal black solid, dashed and dotted lines for each $M_{1}$. We find from the figure that $Y_{\eta_{R}}$ leaves $Y_{\eta_{R}}^{\rm eq}$ around $z_{2}=20$ in these cases. It is considered to be caused due to the break of the balance between the $\eta_{R}$ production by the $N_{2}$ decay and its disappearance due to both the decay and the coannihilation. It shows that the late time additional $\eta_{R}$ production through the $N_{2}$ decay affects the estimation of the $N_{1}$ relic abundance. It forces us to change the strength of the $\eta_{R}$ decay and the $\eta_{R}$ coannihilation from the ones in Fig. 1 in order to realize the required $N_{1}$ relic abundancegggSince the $N_{2}$ decay is relevant to the leptogenesis, we do not change the parameters relevant to it here. since we impose on the model the simultaneous explanation of the DM abundance and the baryon number asymmetry.
Since $h_{1}$ does not affect the values of $h_{2,3}$ determined by the neutrino oscillation data as long as $h_{1}$ is small enough, it can be tuned to result in the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$. By tuning $h_{1}$ suitably to $h_{1}=2.3\times 10^{-9}$ and $2.3\times 10^{-8}$ for the case $M_{1}=$150 GeV and 1.5 GeV respectively, $Y_{N_{1}}$ is found to converge to the required $Y_{N_{1}}^{\infty}$ at $z_{2}$ where $Y_{N_{1}}=Y_{N_{2}}$ is realized. These values of $h_{1}$ are much smaller than the ones used in Fig. 1. It is considered to be caused to reduce the effect of $\eta_{R}$ which is additionally produced through the $N_{2}$ decay by making the relative strength of the $\eta_{R}$ coannihilation to the $\eta_{R}$ decay larger. The similar feature can be observed in the evolution of the number density of $N_{1}$ with the keV scale mass also. For example, the DM relic abundance requires $h_{1}=6.5\times 10^{-5}$ for $M_{1}=5$ keV, which can be confirmed not to affect the result of leptogenesis. In the case $M_{1}=1.5$ TeV, we find that the relation $Y_{N_{1}}+Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ cannot be changed by tuning the value of $h_{1}$. This can be understood if we note $Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ at $z_{2}=300$ where the $\eta_{R}$ coannihilation is frozen out. It suggests that the freezeout of the $\eta_{R}$ coannihilation plays an important role and then $\lambda_{3}$ and $\lambda_{4}$ have to be tuned when $M_{1}$ takes a similar order value to $M_{\eta_{R}}$.
In the right panel, $Y_{N_{1}}$ and $Y_{\eta_{R}}$ are plotted for the case where $\lambda_{3}$ and $\lambda_{4}$ are shifted from the ones in the left panel to $(\lambda_{3},\lambda_{4})=(2.5,-1.0)$ keeping other parameters unchanged. While $Y_{N_{1}}$ in the case of $M_{1}=150$ GeV and 1.5 GeV is found to take a little bit smaller value than the one in the left panel, $Y_{N_{1}}+Y_{\eta_{R}}$ converges to $Y_{N_{1}}^{\infty}$ at $z_{2}=300$ in the case $M_{1}=1.5$ TeV. Since the $\eta_{R}$ coannihilation becomes stronger due to a larger value of $\lambda_{3}$, the contribution from $\eta_{R}$ is considered to be reduced.
@@
These examples show that $N_{1}$ in a wide mass range can explain the required relic abundance of DM in a consistent way with successful leptogenesis by tuning $h_{1}$ and $\lambda_{3,4}$ suitably. Couplings $\lambda_{3}$ and $\lambda_{4}$ play a crucial role especially when $M_{1}$ and $M_{\eta_{R}}$ take the same order values since $Y_{N_{1}}^{\infty}~{{}^{>}_{\sim}}~Y_{\eta_{R}}^{\infty}$ should be satisfied in that case. The model presents an interesting possibility such that three unsolved problems in the SM can be explained by low scale right-handed neutrinos simultaneously.
@@
Finally, we should address phenomenological signatures of the model. Extra charged and neutral scalars $\eta^{\pm}$ and $\eta_{R,I}$ might be directly examined in the future collider experiments [35]. However, since $N_{1}$ has a very small $h_{1}$, it seems to be difficult to distinguish the model from the one in which DM is $\eta_{R}$ through collider signatures. The existence of right-handed neutrinos and the additional inert doublet scalar at TeV scales are expected to cause a new non-negligible contribution at one-loop level to LFV processes like $\mu\rightarrow e\gamma$ and also the electric dipole moment of electron (EDME). It is useful to examine whether their present experimental bounds are satisfied and also to give their prospect in the future experiments. If we use the neutrino Yukawa couplings fixed by eqs. (7) and (12), we can estimate them definitely. Branching ratio of the LFV process $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ in the scotogenic model is expressed as [36]
@@
$B(\ell_{\alpha}\rightarrow\ell_{\beta}\gamma)\simeq\frac{3\alpha}{64\pi(G_{F}M_{\eta}^{2})^{2}}\left|\sum_{k=1}^{3}U_{\alpha k}^{\ast}U_{\beta k}h_{k}^{2}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\right|^{2},$ (28)
@@
$d_{e}/e\simeq\sum_{k=1}^{3}\frac{U_{ek}^{\ast}U_{ek}h_{k}^{2}m_{e}}{16\pi^{2}M_{\eta}^{2}}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\sin\gamma_{k},$ (29)
@@ 4 Summary
The existence of both neutrino mass and DM is unexplained problem in the SM. If it is solved in a common framework, it could be a promising candidate for new physics. The scotogenic model is such an example. In this study, we propose a scenario in the original scotogenic model which allows the lightest right-handed neutrino to be DM without introducing extra interactions. Since this DM cannot be detected through direct detection experiments, such a model may be favorable from the present experimental status.
We discuss remarkable features of the active neutrino mass matrix in the model and show that the right-handed neutrino with mass in a wide rage from TeV to keV can explain the required DM abundance. The model can also explain the baryon number asymmetry through leptogenesis based on the decay of the right-handed neutrino if we assume mass degeneracy among $Z_{2}$ odd ingredients. Although the required mass degeneracy is strict, it is noticeable that three unsolved problems in the SM can be explained through the right-handed neutrinos simultaneously. An interesting point of this scenario is that the relic abundance of DM is determined not only by the interaction of DM itself but also by the interaction of its mother particle $\eta_{R}$. The model gives an example that the interaction searched by the direct detection experiments of DM is irrelevant to the interaction which determines the DM relic abundance.